# Language in the Flow of Time: Time-Series-Paired Texts Weaved into a Unified Temporal Narrative

### v2

@@ E.7 Full Ablation Study Results Using Different Text Encoders
Figure 13: Performance comparison of different text encoders within the TaTS framework. Specifically, we evaluate BERT-110M, GPT2-1.5B, and LLaMA2-7B across multiple datasets using PatchTST (transformer-based model), DLinear (linear-based model), and FiLM (frequency-based model). The results demonstrate that TaTS maintains relatively stable performance across various models and datasets.
@@
Figure 14: Efficiency and forecasting performance comparison of uni-modal time-series modeling, MM-TSFLib, and our TaTS framework. While TaTS incurs a slight increase in training time due to the augmentation of auxiliary variables projected from paired texts, it significantly enhances forecasting accuracy, achieving a lower MSE.

### v3

@@ E.7 Full Ablation Study Results Using Different Text Encoders
Figure 13: Performance comparison of different text encoders within the TaTS framework. Specifically, we evaluate BERT-110M, GPT2-1.5B, and LLaMA2-7B across multiple datasets using PatchTST (transformer-based model), DLinear (linear-based model), and FiLM (frequency-based model). TaTS maintains relatively stable performance across various models and datasets.
In the main experiments, the text embeddings are obtained by a mean pooling over token embeddings. There are more advanced embedding techniques. To this end, we conducted an additional experiment using LLaMA-3.2-1B to generate sentence-level embeddings, while keeping all other settings the same. The results are shown below in Table 30.
Table 30: Mean Square Error of TaTS forecasting with different text embedding methods.
Text Embedding Method Agriculture Climate Economy Security Social Good Traffic
GPT-2 average pooling (prediction length 6) 0.067 1.020 0.008 107.113 0.942 0.174
LLaMA-3.2-1B sentence embeddings (prediction length 6) 0.067 1.015 0.008 106.625 0.935 0.179
GPT-2 average pooling (prediction length 12) 0.153 1.033 0.008 114.754 1.045 0.213
LLaMA-3.2-1B sentence embeddings (prediction length 12) 0.152 1.021 0.008 114.219 1.025 0.209
@@
Figure 14: Efficiency and forecasting performance comparison of uni-modal time-series modeling, $M$-TSFLib, and our TaTS framework. While TaTS incurs a slight increase in training time due to the augmentation of auxiliary variables projected from paired texts, it significantly enhances forecasting accuracy, achieving a lower MSE.